A rectangular lot measures 66m by 107m A sidewalk surrounds the lot The sidewalk is 2m wide what would the area of the sidewalk be
Answer by ankor@dixie-net.com(22740)

A rectangular lot measures 66m by 107m A sidewalk surrounds the lot The sidewalk is 2m wide what would the area of the sidewalk be
:
The overall dimensions including the surrounding sidewalk: 70 by 111 (add 4 m)
:
Overall area - lot area = sidewalk area
(70*111) - (66*107) =
7770 - 7072 = 708 sq/m is the sidewalk area